OpenStudy (anonymous):

postman sold some 15 cent stamps and some 25 cent stamps. 10 stamps total were sold for $1.70. How many of each were sold?

OpenStudy (radar):

Let x = the number of 15 cent stamps sold then 10 - x = number of 25 cent stamp sold The total sold came to $1.70 15x + 25(10-x)=1.70 15x+250-25x=1.70 -10x=170-250-30 -10x=-80 x=8 15 cent stamps 10-8=2 25 cent stamps. $1.20 + .50 =$1.70

OpenStudy (radar):

typo * -10x=170-250=80